Southeastern Bell stocks a certain switch connector at its central warehouse for supplying field service offices. The yearly demand for these connectors is 15,200 units. Southeastern estimates its annual holding cost for this item to be ​$22 per unit. The cost to place and process an order from the supplier is ​$72. The company operates 300 days per​ year, and the lead time to receive an order from the supplier is 2 working days.

​a) What is the economic order​ quantity? units ​(round your response to the nearest whole​ number).

​b) What are the annual holding​ costs? ​$   ​(round your response to the nearest whole​number).

​c) What are the annual ordering​ costs? ​$ ​(round your response to the nearest whole​ number).

​d) What is the reorder​ point? units ​(round your response to the nearest whole​ number).
Answer #1

A. EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST)

EOQ = SQRT(2 * 15200 * 72 / 22) = 315.42

B. ANNUAL HOLDING COST = (EOQ / 2) * HOLDING COST

ANNUAL HOLDING COST = (315.42 / 2) * 22 = EOQ = SQRT(2 * 15200 * 72 / 22) = 315.42

C. ANNUAL ORDERING COST = (DEMAND / EOQ) * ORDERING COST

ANNUAL ORDERING COST = (15200 / 315.42) * 72 = 3469.66

D. DAILY DEMAND = 15200 / 300 = 50.67

ROP = DAILY DEMAND * LEAD TIME

= 50.67 * 2 = 101.34